Suppose that a researcher studying the weight of female college athletes wants to predict the weights based on height, measured in inches, the percentage of body fat of an athlete, and age. The researcher calculates the regression equation as (weight) = 4.767*(height) + 0.687*(percent body fat) - 1.028*(age) - 85.889. If a female athlete is 66 inches tall, has a 17 percentage of body fat, and is 21 years old, what is her expected weight?
Answer #1

from given regression equation: (weight) = 4.767*(height) + 0.687*(percent body fat) - 1.028*(age) - 85.889 :

expected weight =4.767*66+0.687*17-1.028*21-85.889=218.824